OpenStudy (anonymous):

When a 3-kg iron skillet is heated on the stove, it becomes hot very quickly. When 3 kg of water is heated on the stove it heats up slowly. What is the reason for this difference? Support your answer using 3 – 4 complete sentences.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@KyanTheDoodle

OpenStudy (anonymous):

@cramos725

OpenStudy (michele_laino):

let's suppose to heat the same mass of water and iron, namely M=3 Kg of iron and M=3 kg of water Let's suppose furthermore, that the temperature difference is the same, say \[\Delta \theta \]

OpenStudy (michele_laino):

then the quantity of heat absobed by iron is: \[\Delta {Q_{IRON}} = {c_{IRON}}M\Delta \theta \]

OpenStudy (michele_laino):

where the quantity of heat absorbed by water is: \[\Delta {Q_{WATER}} = {c_{WATER}}M\Delta \theta \]

OpenStudy (anonymous):

None of that makes sense to me

OpenStudy (michele_laino):

now we can consider the specific heat of a substance: \[\Large c = \frac{1}{M}\frac{{\Delta Q}}{{\Delta \theta }}\] as a heating rate with respect to the temperature difference

OpenStudy (michele_laino):

empirically we can say that the subsequent condition holds: \[\Large \Delta {Q_{IRON}} > \Delta {Q_{WATER}}\] from which, we get: \[\Large {c_{IRON}} > {c_{WATER}}\]

OpenStudy (michele_laino):

and keeping in mind the meaning of c as heating rate, we get your answer

OpenStudy (michele_laino):

another reasoning is: the mobility of electrons inside the iron (please think about the Fermi sea) is greater than the mobility of the molecule of water, so the heat will transmit itself more rapidly in iron than in water